OpenStudy (anonymous):

Simplify i^22 1 –1 –i i

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

\[i^{22} = {(i^{2})}^{11} = (-1)^{11} = -1\]

OpenStudy (angela210793):

(i^2)^11=(-1)^11=-1

OpenStudy (anonymous):

I prefer this one: \[i^2 = -1 \implies i^4 = (-1)^2 = 1\]\[i^{22} = i^{4\cdot 5}\cdot i^2= (i^4)^5 \cdot i^2 = 1^5\cdot -1 = -1\]